Dr. Oats invests $100,000 in a piece of land that is expected to increase in value by 12% per year for the next five years. She will then take the proceeds and provide herself with a 10-year annuity. Assuming a 14% interest rate for the annuity, how much will this annuity be?
Answer #1


solution: FV=PV*(1+r)^n FV=100000*(1+0.12)^5 FV=Value of land at the end of 5 years **factor working : PVIFA(R%,n) PVIFA(14%,